OpenStudy (anonymous):

y = -x + 2; 7x + 4y = -1

OpenStudy (texaschic101):

y = -x + 2 --- now sub -x + 2 in for y in the other equation 7x + 4y = -1 7x + 4(-x + 2) = -1 --- distribute through the parenthesis 7x - 4x + 8 = -1 --- subtract 8 from both sides 7x - 4x = -1 - 8 -- combine like terms 3x = - 9 -- divide both sides by 3 x = -3 now sub -3 in for x in either equation to find y y = -x + 2 y = - (-3) + 2 y = 3 + 2 y = 5 check.. 7x + 4y = -1 7(-3) + 4(5) = -1 -21 + 20 = -1 -1 = -1 (correct) x = -3 and y = 5